Last year Miami Rivets had S6 million in operating income (EBIT). Its depreciation expense was $1 million, its interest expense was $1 min, and its corporate tax rate was 30%. At year-end, it had $15 million in current assts, $4 million in accounts payable, $1 million in accruals, $2 million in notes payable, and $16 million in net plant and equipment. Miami Rivets uses only debt and common equity to fund its operations. (In other words, Miami Rivets has no preferred stock on its balance sheet.) Miami Rivets had no other current liabilities. Assume that Miami Rivetss only noncash item was depreciation A. What was the companys net income? B. What was its net operating working capital (NowC)? C. What was its net working capital (NWC)? CWhat was its net working capital (NWC? D. Miami Rivets had $12 million in net plant and equipment the prior year. Its net operating working capital has remained constant over time. What is the companys free cash flow (FCF) for the year that just ended?
Answer #1
1)
A) EBIT 6000000
Interest 1000000
EBT 5000000
Tax at 30% 1500000
Net income 3500000
B) Current assets 15000000
ess:
Accounts payable 4000000
Accruals 1000000
Net operating working capital 10000000
C) Net working capital = Current assets-Current liabilities = 15000000-4000000-1000000-2000000 = 8000000
D) EBIT 6000000
Tax at 30% 1800000
NOPAT 4200000
Add: Depreciation 1000000
Operating cash flow 5200000
Less:
Net capital spending = Ending net plant and equipment-Beginning net plant and equipment+Depreciation expense = 16000000-12000000+1000000 = 5000000
Change in NWC 0
Free cash flow 200000
E) Common stock 5000000
Retained earnings ending balance:
Beginning balance 11200000
Net income for the year 3500000
Dividends paid -1200000
Retained earnings ending balance 13500000
EOY stockholders' balance 18500000
F) MVA = MV of stock-BV of stockholders' equity
= 50*500000-18500000 = 6500000
G) EVA = NOPAT-Invested capital*WACC
= NOPAT-(NOWC+NFA)*WACC = 4200000-(10000000+16000000)*9% = 1860000
2) Tax table not given